OpenStudy (anonymous):

how do i find this answer.... 215.4% of 55= 118.47... I have a TEAS test tomorrow and i cant figure out how to find this answer!

OpenStudy (hoblos):

55* 2.154 = 118.47

OpenStudy (anonymous):

'per cent' means out of 100 so 215.4% is like saying 215.4/100 so 215.4/100 x 55=118.47

OpenStudy (anonymous):

you convert the percent into a decimal and multiply it to the number 215.4% = 2.154 2.154 * 55 = 118.47

OpenStudy (anonymous):

thank you for all of your help!

OpenStudy (anonymous):

ok what about a word problem that is basically percent has gone up from 625% to 710% in one year... what is the percent cent of increase to the nearest tenth

OpenStudy (anonymous):

625 to 710 over one year what is the % increase to the nearest tenth

OpenStudy (anonymous):

625 to 710 over one year what is the % increase to the nearest tenth

OpenStudy (anonymous):

if you mean for the answer to be 710 - 625 then there is no tenth, and by asking for the increase, that would be how you find it, so I'm confused as to what your asking.

OpenStudy (anonymous):

the enroat a college went up from 625 710 students over a period of 1 year. what is the percent increase to the nearest tenth? ... thats the question

OpenStudy (anonymous):

for that question, you would find the difference between the two (710-625 = 85) then find what percent that is compared to the original number of students which is 625 (do so by dividing 85 by 625 which should be something close to .13, you should check that) whatever that number is times 100 should be the % increase. (i think it should be around 13% but you should check my math because i don't have a calculator with me.